How do you complete the square for the function: h=-4.9t^2+24.5t+6
Answer by richwmiller(17219)

We can factor it!
h=-4.9(t-5.23395)(t+0.233951)
h=-4.9*(t^2-5t-1.22449)
t^2-5t
take 5/2 and square it
25/4
h+25/4+4.9(1.22449)=-4.9(t^2-5t+25/4)